OpenStudy (lorenbeech):

If anyone has the time and patience to help me please do :) The question is below:

OpenStudy (lorenbeech):

1.) Let C(t) be the number of cougars on an island at time t years (where t > 0). The number of cougars is increasing at a rate directly proportional to 3500 . C(t). Also, C(0) = 1000, and C(5) = 2000. A. Find C(t) as a function of t only. B. Calculate C(10). C. Find the limit as t tends to infinity of C(t) , and explain its meaning.

OpenStudy (lorenbeech):

For A I got C(t) = 3500 – 2500 e^((ln .6)/5 * T) bit I'm unsure of myself

OpenStudy (lorenbeech):

but*

OpenStudy (agent0smith):

Show how you got it so we can check?

OpenStudy (agent0smith):

3500 . C(t) not sure what this is... 3500*C(t)? If it is that \[\Large \frac{ dC }{ dt } = 3500 C\]

OpenStudy (lorenbeech):

dC/dt = k ( 3500 – C(t)) dC/dt+ kC(t) = 3500k e^(kt)(dC/dt + kC(t) ) = 3500k e^(kt) (e^(kt)C(t))’ = 3500k e^(kt) integrate e^(kt)C(t) = 3500e^(kt) + A C(t) = 3500 + Ae^(-kt) when t = 0: 1000 = 3500 + A A = -2500 t = 5: 2000 = 3500 – 2500 e^(-5k) 2500e^(-5k) = 1500 e^(-5k)= .6 -5k = ln .6 k = -(ln .6)/5 C(t) = 3500 – 2500 e^((ln .6)/5 * T)

OpenStudy (faiqraees):

Does 3500 . C(t). means 3500 - C(t) ?

OpenStudy (agent0smith):

If you have \[\large \frac{ dC }{ dt } = k ( 3500 – C(t)) \] then it's a separable equation \[\large \frac{ dC }{ 3500-C(t) }=k dt\]

OpenStudy (faiqraees):

Yes this method is much better and fast

OpenStudy (agent0smith):

\[\large \int\limits \frac{ dC }{ 3500-C(t) }= \int\limits k dt\] \[\Large -\ln(3500-C(t)) = kt+B\] I'm using B as the constant of integration, because C is taken. Then just solve for C(t).

OpenStudy (agent0smith):

But your method worked fine.

OpenStudy (faiqraees):

She multiplied the equation by negative 1 afterwards

OpenStudy (faiqraees):

2000 = 3500 – 2500 e^(-5k) -1500 = -2500 e^(-5k) -1 x -1500 = -1 x -2500 e^(-5k) 1500 = 2500e^(-5k) 2500e^(-5k) = 1500

OpenStudy (agent0smith):

B. Calculate C(10). C. Find the limit as t tends to infinity of C(t) , and explain its meaning. These are easy now. B. plug in t=10 C. is a little tougher but not hard... just find \[\large \lim_{t \rightarrow \infty} \left( 3500 – 2500 e^{t* \ln .6/5 }\right)\]

OpenStudy (agent0smith):

Remember that ln(0.6) is negative (because I didn't)

OpenStudy (lorenbeech):

Thank you both

OpenStudy (lorenbeech):

Just to clarify, B. would be 3470?

OpenStudy (lorenbeech):

@agent0smith

OpenStudy (agent0smith):

Idk, I didn't work it out, just plug it in on wolframalpha or something to check, cos I gotta go

OpenStudy (agent0smith):

But no it's not 3470

OpenStudy (lorenbeech):

I'm sorry Calculation error. I got 1025.41125 instead?

OpenStudy (agent0smith):

No... not close... show your work, but i gotta go anyway all you have to do is plug in t=10 into: 3500 – 2500 e^((ln .6)/5 * T) and type it exactly like that into a calculator or w/e

OpenStudy (lorenbeech):

2600??? @agent0smith

OpenStudy (agent0smith):

Yes

OpenStudy (lorenbeech):

Gosh dangit lol thanks. Now for the limit I got 3500. Is that correct?

OpenStudy (agent0smith):

Yes.

OpenStudy (lorenbeech):

Now what does it mean by "explain it's meaning"? Like 3500 is the limit of the amount of cougars?

OpenStudy (lorenbeech):

@agent0smith

OpenStudy (agent0smith):

Exactly!